# Stateful Large Language Model Serving with Pensieve

### v1

@@ Abstract.
Large Language Models (LLMs) have recently experienced great success, as evident in the widespread popularity of ChatGPT. Existing LLM serving systems are stateless across requests. Consequently, when LLMs are used in the common setting of multi-turn conversations, a growing log of the conversation history must be processed alongside any request by the serving system at each turn, resulting in repeated history processing.
In this paper, we design $Pensieve$, a system optimized for multi-turn conversation LLM serving. $Pensieve$ maintains the conversation state across requests by caching previously processed history to avoid duplicate processing. $Pensieve$’s multi-tier caching strategy can utilize both GPU and CPU memory to efficiently store and retrieve cached data. $Pensieve$ also generalizes the recent PagedAttention kernel to support attention between multiple input tokens with a GPU cache spread over non-contiguous memory. Our evaluation shows that Pensieve is able to achieve 1.51-1.95x throughput compared to vLLM and reduce latency by 60-75%.
@@ 1. Introduction
The world has recently witnessed the fast expansion of Large Language Models (LLMs). The most popular use of LLMs are for chatbots, with applications like ChatGPT demonstrating an astounding capability of chatting and answering questions like real humans. Other LLM-backed applications are also being applied to automate a variety of tasks, such as writing ads, responding to emails, doing literature review etc. As LLM continues its explosive growth, it is imperative to develop fast and efficient LLM serving systems.
@@
An LLM is an autoregressive DNN model based on Transformer (transformer, ) architecture. The model iteratively predicts the next output token based on the current context which includes the sequence of input prompt tokens followed by output tokens generated in the previous iterations. LLMs require very expensive computation due to two factors. One, LLMs have huge parameter sizes (10s or 100s of billions) and have a trend of growing even larger. Two, LLMs need to support a large context size (2K to 32K tokens) to be useful. There has been much related work which improve the performance of LLM inference/serving from various angles, including better batching (orca, ), operation fusion (flash, ), better GPU memory utilization (vllm, ), faster output generation (speculative:leviathan, ; speculative:chen, ), low rank adaptation (lora, ) and quantization (qlora, ) (see §8). In contrast to these work, in this paper, we take a step back and examine inefficiencies that arise in the context of a specific but very popular LLM use case today, aka as a multi-turn conversational chatbot.
@@
In the conversational setup, the user and the chatbot are engaged in a dialogue that may last many rounds. In order for the chatbot not to “lose memory” of what has been said so far when responding, the cumulative history of the dialogue must be part of the context for LLM’s autoregressive generation. As existing LLM serving systems are stateless across requests, clients send a growing log of conversation history alongside each new request as the input prompt to be processed from scratch. This causes much duplicate processing for multi-turn conversations.
@@
How to avoid duplicate processing of the chat history? To do so, the serving system can save any previously processed context data. When new requests from the same conversation arrive, the saved context data can be re-used and subsequently augmented. This can be done with best effort. Essentially, the serving system is allowed to keep some cached state containing processed context across requests. Doing so enables the serving system to exploit the opportunity that when users are actively chatting with an AI chatbot, follow-up requests usually arrive within a reasonably short time period to leverage cached state.
@@
Caching state across requests is straightforward at the high-level, but several challenges remain to make it really work. First, where to save the data? Keeping it in the GPU is the fastest, but is very constrained by the relatively small GPU memory size. Putting it on disk would incur much longer load latency, hurting the user experience. A two-tier caching solution spanning both GPU and CPU memory is promising, but care must be taken to cope with each tier’s capacity limit and to swap in/out the saved context data from/to the GPU without damaging performance. Second, how to re-use saved context data efficiently when processing a new request? When restoring a conversation’s context data from CPU to GPU, how to merge it with the conversation’s existing context that is already cached in the GPU? Furthermore, what has been saved might not be complete due to the cache limit. How to handle dropped cache?
@@
In this paper, we design a stateful LLM serving system, called Pensieve, to address the aforementioned challenges. Pensieve saves a conversation’s processed context in a two-tier GPU-CPU cache. It evicts cached data to the next tier (or discards it) according to how long the conversation has been inactive and how large the conversation’s saved context has accumulated. The eviction is done at the granularity of a suffix of the conversation instead of the whole. Therefore, a conversation’s saved context might span both tiers of the cache and also be partially trimmed. Pensieve uses ahead-of-time swapping and pipelined transfer to overlap computation with the data movement between cache tiers. Dropped contexts are handled via recomputation. Evicting and restoring cause a conversation’s cached context to occupy non-contiguous GPU memory. We develop a new GPU kernel to compute attention (transformer, ) between multiple input tokens and cached context residing in non-contiguous memory, which lacks in existing LLM serving systems. Our kernel is a generalized version of the PagedAttention kernel in vLLM (vllm, ).
@@
We have built the Pensieve and compared its performance against vLLM (vllm, ), a state-of-the-art serving system that does not cache state across requests. Experiments show that Pensieve is able to improve serving throughput by 1.51-1.95x and reduce normalized latency by 60-75%.
@@
  • We build a new attention GPU kernel to compute attention between a new request’s multiple input tokens and the saved context scattered in non-contiguous GPU memory. Existing kernels either require contiguous GPU context cache or restrict to a single input token.
  • We evaluate Pensieve using real world conversation datasets to demonstrate its effectiveness compared to the state-of-the-art stateless serving system.
@@ 2. Background
Figure 1. Inference of Transformer-based Large Language Models
In this section, we provide a brief background on LLM, its core attention operation, and how existing systems serve LLMs.
@@ 2.1. LLM and the Attention Mechanism
Popular large language models, e.g. GPT-3 (gpt3, ), OPT (opt, ), Llama (llama, ), are all based on the Transformer architecture (transformer, ). A model consists of many transformer layers, each of which is composed of an attention module, seen as the dashed box Figure 1, and a 2-layer feed forward network. The model takes as input a sequence of token ids representing the natural language sentence and feeds them through an embedding layer to obtain a continuous representation (aka embedding) for each token before feeding them through transformer layers. For simplicity, we refer to token embeddings as “tokens”, and refer to token ids as “raw tokens”. LLM is autoregressive in the sense that it iteratively predicts the next output token based on the current context which includes the input prompt tokens followed by output tokens generated in the previous iterations.
The success of Transformer originates from the capability of its attention module. For each layer, the attention module first performs linear transformations on its input token embedding ($X$) to produce three new embeddings Query ($Q$), Key ($K$), and Value ($V$):
@@
The above equations show the process of a so-called single-head attention. In practice, multi-head attention is used, where $Q$, $K$, $V$ produced by Equation (1) are divided into groups, called attention heads. Each attention head independently performs attention as Equation (2) and (3).
@@ 2.2. How LLM is Served
To perform inference using an LLM, one needs to keep a KV cache in GPU to avoid recomputation during the autoregressive output generation. Figure 1 shows the typical LLM inference process adopted by systems like FasterTransformer (fastertransformer, ), ORCA (orca, ), vLLM (vllm, ). It is divided into two phases: 1) The prefill phase processes all the input prompt tokens all together to generate the $K$ and $V$ token embeddings (aka KV-tokens) for each Transformer layer and initialize the KV cache with the resulting KV-tokens. The embedding of the last token of the prompt from the last layer’s $O$ tensor is used to generate the first output token; 2) The iterative generation phase takes in the token generated by last step as a single new input token. For each layer, it computes the $Q$,$K$,$V$ embedding for the new token, updates the KV cache, and performs attention using the new token’s $Q$ embedding with the entire context’s KV-tokens.
@@
For LLMs have variable sized input and output, batching granularity has a huge impact on system throughput and serving latency. If scheduling is performed at the request granularity, executing a batch of requests with different input prompt lengths require padding tensors to the maximum length and waiting for the request with the longest output to finish. Iteration-level batching strategy, originally proposed by BatchMaker (batchmaker, ) for non-transformer-based sequence-to-sequence models, performs batching at token granularity. Whenever a request finishes an iterative generation step, the scheduler checks whether it has reached the end of sequence and can leave the batch, making room for a request to start its computation immediately. As we advance into Transformer era, ORCA (orca, ) extends this approach to support the LLM workload.
@@
For each request, the model performs iterative generation until either the special end-of-sentence token (EOS) is emitted or the preconfigured maximum decoding length is reached. Systems like FasterTransformer (fastertransformer, ) and ORCA (orca, ) reserve slots in KV cache for each request based on the maximum decoding size. A more recent system, vLLM (vllm, ), can dynamically grow the allocated cache slots for each request and allow these slots to reside in non-contiguous GPU memory. vLLM develops PagedAttention to handle the generation phase with non-contiguous KV cache. Existing serving systems are stateless across requests. In other words, they de-allocate all the cache slots used by a request as soon as it finishes.
@@ 3.1. Motivation
Figure 2. Existing serving systems send a cumulative history with each request in a multi-turn conversation.
Existing techniques for serving Large Language Model mostly focus on improving the inference time or memory efficiency of a single request. We take a step back and examine inefficiencies that arise in the context of a specific but very popular LLM use case today, aka as a multi-turn conversational chatbot.
In a multi-turn conversation, user has multiple rounds of conversations with the chatbot and LLM needs to be aware of the conversation history in order to generate appropriate response. This is done by prefixing each new request with the entire cumulative conversation history, due to the stateless nature of existing serving systems, as shown in Figure 2. As the interaction between user and chatbot goes on, conversation history grows, making the cost of the prefill phase soon overshadow that of iterative generation phase. Unfortunately, much of the history processing is redundant.
@@
The goal of this project is to minimize redundant computation of the conversation history. This can be done by caching any previously processed embeddings at the serving system and re-use them across requests from the same conversation. More concretely, one can save the KV-tokens in the KV cache belonging to a previous request and only process the new prompts for the next follow-up request.
Figure 3 demonstrates the heavy cost of prompt initiation phase under an artificial workload where each request generates 200 tokens and has varying conversation history sizes. As shown in the figure, the cost of recomputing the conversation history (solid blue line) causes the cost of the prefill phase to soon outgrow the generation phase.
@@ 3.2. Challenges
Limite GPU memory for caching.
@@
LLM has very large model parameters, which results in large KV-tokens. For example, a 13 billion parameter GPT-3 model has 40 layers and a hidden size of 5120. Assuming the use of 16-bit half precision numbers, storing each KV-token takes 2 * 40 (layer) * 5120 (units/layer) * 2 (bytes/unit) = 0.78MB memory space. Under the current GPU memory limit, depending on the history lengths, only a few dozens or hundreds of conversation histories can be kept in the GPU. We must extend our cache space to use the more abundant CPU memory.
Cache management.
When using a multi-tier GPU-CPU cache, the serving system must be able to evict and restore cached data. The system has to make best-effort decisions to swap cached KV-tokens from GPU to CPU to make room for requests from other conversations. Later, it needs to recover the swapped out KV-tokens as fast as possible. When the system is under CPU memory pressure, some cached KV-tokens might need to be dropped. Therefore, we need a low overhead solution to manage memory space and swapping and deal with the case that some parts of the conversation history embeddings are no longer available in the system.
Handling non-contiguous KV cache
@@
Figure 4. Multi-Query Attention with Non-contiguous Context Cache. Batch contains two requests. Left-hand side shows the Query representation for tokens being processed in each request. Right-hand side shows the context data (Key and Value representation) for each request in GPU.
In the multi-turn conversation setup, whenever a new requests comes in an existing conversation, the request’s new prompt, which in all likelihood has more than one input token, needs to be processed along cached KV-tokens as part of the prefill phase. For any cached KV-token that has been previously evicted to the CPU, there is no guarantee they are restored to the GPU in contiguous memory region next to existing KV-tokens cached in the GPU. Figure 4 provides an illustration of the computation pattern of attention. The batch contains request 0 which is currently performing generation using the last generated token, and request 1 which just arrives with a 7-token new prompt. To perform attention operation, request 0 needs to iterate over tokens in KV cache block 3 and 1, and request 1 needs to iterate over tokens in cache block 0, 4, 5, 2, 7. Existing high performance attention kernels, e.g. MemoryEfficientAttention (memeff, ), FlashAttention (flash, ), require inputs to reside in contiguous memory space and therefore cannot work with non-contiguous paged memory management. Although vLLM has developed the PagedAttention GPU kernel, it is designed to be solely used in the generation phase, and thus has the limitation that each request in the batch can have at most one input token. Hence, PagedAttention kernel cannot support prompt initiation for request 1 in Figure 4. And due to the design of single-query PagedAttention, no trivial way is available to extend it to support multi-query attention. An alternative approach might be to process the new prompt query in the same iterative manner as incremental decoding, one token a time, so that PagedAttention kernel can be applied. But this approach gives up the parallelization opportunity available in the prefill phase due to the extra query dimension, and will lead to significantly longer computation time for long input prompts.

### v2

@@ Abstract.
Large Language Models (LLMs) are wildly popular today and it is important to serve them efficiently. Existing LLM serving systems are stateless across requests. Consequently, when LLMs are used in the common setting of multi-turn conversations, a growing log of the conversation history must be processed alongside any request by the serving system at each turn, resulting in repeated processing.
In this paper, we design $Pensieve$, a system optimized for multi-turn conversation LLM serving. $Pensieve$ maintains the conversation state across requests by caching previously processed history to avoid duplicate processing. $Pensieve$’s multi-tier caching strategy can utilize both GPU and CPU memory to efficiently store and retrieve cached data. $Pensieve$ also generalizes the recent PagedAttention kernel to support attention between multiple input tokens with a GPU cache spread over non-contiguous memory. Our evaluation shows that $Pensieve$ can achieve 13-58% more throughput compared to vLLM and TensorRT-LLM and significantly reduce latency.
@@ 1. Introduction
The world has recently witnessed the fast expansion of Large Language Models (LLMs). The most popular use of LLM is for chatbots, with applications like ChatGPT demonstrating an astounding capability of following instructions and interacting with humans as a virtual assistant. Other LLM-backed applications include writing code, responding to emails, doing literature reviews, etc. As LLM continues its explosive growth, it is imperative to develop fast and efficient LLM serving systems.
@@
An LLM is an autoregressive DNN model based on the Transformer (Vaswani et al., 2017) architecture. The model iteratively predicts the next output token based on the current context which includes the sequence of input prompt tokens followed by output tokens generated in the previous iterations. LLMs require very expensive computation due to two factors. One, LLMs have huge parameter sizes (10s or 100s of billions) and have a trend of growing even larger. Two, LLMs need to support a large context size (2K to 32K tokens) to be useful. There has been much related work to improve the performance of LLM inference/serving from various angles, including better batching (Yu et al., 2022), operation fusion (Dao et al., 2022), better GPU memory utilization (Kwon et al., 2023), faster output generation (Leviathan et al., 2023; Chen et al., 2023), low rank adaptation (Hu et al., 2021) and quantization (Dettmers et al., 2023) (see §7). In contrast to these works, in this paper, we take a step back and examine inefficiencies that arise in the setting of a specific but very popular LLM use case today, aka as a multi-turn conversational chatbot.
@@
In the conversational setup, the user and the chatbot are engaged in a dialogue that may last many rounds. In order for the chatbot not to “lose memory” of what has been said so far when responding, the cumulative history of the dialogue must be part of the context for LLM’s autoregressive generation. As existing LLM serving systems are stateless across requests, one must prepend a growing log of conversation history alongside each new request as the input prompt to be processed from scratch. This causes much duplicate processing for multi-turn conversations.
@@
How to avoid duplicate processing of the chat history? To do so, the serving system can save any previously processed context data in the form of token embeddings. When new requests from the same conversation arrive, the saved context data can be re-used and subsequently augmented. This can be done with best effort. Essentially, the serving system is allowed to keep some cached state containing previously processed context across requests. Doing so enables the serving system to exploit the opportunity that when users are actively chatting with an AI chatbot, follow-up requests usually arrive within a reasonably short time period to leverage the cached state.
@@
Caching state across requests is straightforward at the high level, but several challenges remain to make it really work. First, where to save the data? Keeping it in the GPU is the fastest, but is very constrained by the relatively small GPU memory size. Putting it on disk would incur much longer load latency, hurting the user experience. A two-tier caching solution spanning both GPU and CPU memory is promising, but care must be taken to cope with each tier’s capacity limit and to swap in/out saved context data from/to the GPU without damaging performance. Second, how to reuse saved context data efficiently when processing a new request? Furthermore, some parts of the saved context can be dropped due to the cache limit. How to handle partially saved context by recomputing what has been dropped?
@@
In this paper, we design a stateful LLM serving system, called Pensieve, to address the aforementioned challenges. Pensieve saves a conversation’s processed context in a two-tier GPU-CPU cache. It evicts cached data to the next tier (or discards it), preferring conversations that have been inactive for longer and/or those that are cheaper to recompute. The eviction is done at the granularity of a chunk of tokens instead of the whole conversation. Therefore, a conversation’s saved context might span both tiers of the cache and may be partially dropped. Pensieve uses ahead-of-time swapping and pipelined transfer to overlap computation with the data movement between cache tiers. Dropped contexts are handled via recomputation. Evicting and restoring cause a conversation’s cached context to occupy non-contiguous GPU memory. We develop a new GPU kernel to compute attention (Vaswani et al., 2017) between multiple input tokens and cached context residing in non-contiguous memory, which is lacking in existing LLM serving systems. Our kernel is a generalized version of the PagedAttention kernel in vLLM (Kwon et al., 2023).
@@
We have built Pensieve and compared its performance against vLLM (Kwon et al., 2023) and TensorRT-LLM (NVIDIA, 2023), state-of-the-art serving systems that do not cache state across requests. Experiments show that Pensieve can improve serving throughput by 13% to 58% and also significantly reduce latency at moderate load.
@@
  • We build a new attention GPU kernel to compute attention between a new request’s multiple input tokens and the saved context scattered in non-contiguous GPU memory. Existing kernels either require contiguous GPU context cache or are restricted to a single input token.
  • We evaluate Pensieve using real-world conversation datasets to demonstrate its effectiveness compared to the state-of-the-art stateless serving system.
@@ 2. Background
Figure 1. How inference is done for Transformer-based LLMs
This section provides a brief background on LLMs and how existing systems serve them.
@@ 2.1. LLM and the Attention Mechanism
Popular large language models, e.g. GPT-3 (Brown et al., 2020), OPT (Zhang et al., 2022), Llama (Touvron et al., 2023a, b), are all based on the Transformer architecture (Vaswani et al., 2017). A model consists of many transformer layers, each of which is composed of an attention module, seen as the dashed box Figure 1, and a 2-layer feed-forward network. The model takes as input a sequence of token IDs representing the natural language sentence and feeds them through an embedding layer to obtain a continuous representation (aka embedding) for each token before feeding them through transformer layers. For simplicity, we refer to token embeddings as “tokens”, and refer to token IDs as “raw tokens”. LLM is autoregressive in the sense that it iteratively predicts the next output token based on the current context which includes the input prompt tokens followed by output tokens generated in the previous iterations.
The success of the Transformer originates from the capability of its attention module. For each layer, the attention module first performs QKV projection, aka linear transformations on its input token embedding ($X$), to produce three new embeddings Query ($Q$), Key ($K$), and Value ($V$):
@@
The above equations show the process of so-called single-head attention. In practice, multi-head attention is used, where $Q$, $K$, $V$ produced by Equation (1) are divided into groups, called attention heads. Each attention head independently performs attention as Equation (2) and (3).
@@ 2.2. How LLM is Served
To perform inference using an LLM, one needs to keep a KV cache in GPU to avoid recomputation during the autoregressive output generation. Figure 1 shows the typical LLM inference process adopted by systems like FasterTransformer (NVIDIA, 2021), ORCA (Yu et al., 2022), vLLM (Kwon et al., 2023). It is divided into two phases: 1) In the prefill phase, all input prompt tokens are processed together to generate $K$ and $V$ (aka KV-tokens) for each layer, and the KV cache is initialized with the resulting KV-tokens. The embedding of the last token from the last layer is used to generate the first output token; 2) The generation (also referred to as decoding) phase works iteratively over many steps. In each step, the token generated by the last generation step is processed as a single new input token. Each layer computes the $Q$,$K$,$V$ embedding vector for the new token, updates the KV cache, and performs attention using the new token’s $Q$ embedding with all KV-tokens in the KV cache.
@@
For LLMs that have variable-sized input and output, the granularity of batching has a huge impact on system throughput and serving latency. If scheduling is performed at the request granularity, executing a batch of requests with different input prompt lengths requires padding tensors to the maximum length and waiting for the request with the longest output to finish. Iteration-level batching strategy, originally proposed by BatchMaker (Gao et al., 2018) for non-transformer-based sequence-to-sequence models, performs batching at token granularity. ORCA (Yu et al., 2022) extends this approach to support the LLM workload: whenever a request finishes an iterative generation step, the scheduler checks whether it has reached the end of a sequence and can leave the batch, making room for a request to start its generation phase immediately.
@@
For each request, the model performs iterative generation until either the special end-of-sentence token (EOS) is emitted or the preconfigured maximum decoding length is reached. Systems like FasterTransformer (NVIDIA, 2021) and ORCA (Yu et al., 2022) reserve slots in KV cache for each request based on the maximum decoding size. A more recent system, vLLM (Kwon et al., 2023), can dynamically grow the allocated cache slots for each request and allow these slots to reside in non-contiguous GPU memory. vLLM develops PagedAttention GPU kernel to handle the generation phase with non-contiguous KV cache. Existing serving systems are stateless across requests. In other words, they de-allocate all the cache slots used by a request as soon as it finishes.
@@ 3.1. Motivation
Figure 2. Existing serving systems process a cumulative history repeatedly with each request in a multi-turn conversation.
Existing techniques for serving LLMs mostly focus on improving the inference time or memory efficiency of a single request. We take a step back and examine inefficiencies that arise in a very popular LLM use case today, aka as a multi-turn conversational chatbot.
In a multi-turn conversation, the user engages in multiple rounds of conversations with the chatbot so the underlying LLM needs to be aware of the conversation history to generate an appropriate response. This is done by prepending the cumulative conversation history as raw text to each new request, due to the stateless nature of existing serving systems, as shown in Figure 2. As the interaction between the user and chatbot continues, the conversation history grows, making the cost of the prefill phase overshadow that of the iterative generation phase. Unfortunately, much of the history processing is redundant.
Figure 3 demonstrates the heavy cost of the prompt initiation phase under an artificial workload where each request has 200 new prompt tokens and has varying conversation history sizes. As shown in the figure, the cost of recomputing the conversation history (solid blue line) causes the cost of the prefill phase to soon outgrow the generation phase.
The goal of this project is to minimize redundant computation of the conversation history. This can be done by caching any previously processed embeddings at the serving system and re-using them across requests from the same conversation. More concretely, one can save the KV-tokens in the KV cache belonging to a previous request and only process the new user prompts of the next follow-up request while re-using history embeddings saved in the KV cache.
@@ 3.2. Challenges
Limited GPU memory for caching.
@@
LLM has very large model parameters, resulting in large KV-tokens. For example, a 13 billion parameter GPT-3 model has 40 layers and a hidden size of 5120. Assuming the use of 16-bit half-precision numbers, storing each KV-token takes 2 * 40 (layer) * 5120 (units/layer) * 2 (bytes/unit) = 0.78MB memory space. Given the limited GPU memory capacity, depending on the history lengths, only a few dozen or hundreds of conversation histories can be kept in the GPU. Therefore, we must extend our cache space to use the more abundant CPU memory.
Token-level cache management and recovery
When using a multi-tier GPU-CPU cache, the serving system needs to swap cached history from GPU to CPU and vice versa. Swapping at the coarse granularity of an entire conversation history is sub-optimal; not only does it utilize the cache space inefficiently but also it incurs large swapping latency. Thus, we decide to swap at the granularity of individual tokens. Specifically, in order to make room for the processing of new requests, the serving system chooses certain cached KV-tokens to swap from GPU to CPU. Later, it also needs to restore the swapped out KV-tokens from CPU to GPU.
Figure 4. Execution time of attention operation for a chunk of 32 tokens with different context sizes. Results are normalized by the execution time of non-attention operations in a transformer layer.
Figure 5. Layout of a typical request’s KV-token context. The shaded areas, which occur at both ends of the context, mark those tokens that must be processed by the prefill phase.
When the system is under CPU memory pressure, some cached KV-tokens need to be dropped and re-computed later when needed. We note that, although each token occupies the same amount of memory space, the recomputation cost of each token is different due to the nature of causal attention computation. Specifically, tokens appearing later in the context sequence require more computation than earlier ones because tokens “attend to” all preceding but not succeeding tokens. Figure 4 shows the execution time of the attention operator for 32 tokens with a prompt context of varying sizes; the execution time shown has been normalized against the rest of inference time (aka the sum of execution times for all other non-attention operators). As can be seen in Figure 4, the cost of attention grows linearly with context size. Thus, when deciding which tokens to discard to reclaim CPU memory, it is more preferable to drop the leading tokens of a conversation history.
Dropping KV-tokens from the leading end of a conversation brings additional complexities. Figure 5 illustrates the layout of a typical request from a continuing conversation in its prefill phase. The request’s context can be viewed as composed of four segments: 1) the first and earliest segment corresponds to tokens that have been dropped from the CPU cache and must be recomputed. 2) the second segment corresponds to tokens that reside in the CPU cache and will be fetched into the GPU. 3) the third segment contains tokens residing in the GPU cache. 4) the fourth and latest segment contains the raw tokens corresponding to this request’s new prompt. As we can see, both the first and fourth segment requires computation. However, such separation of computation at both ends of the context breaks the assumption of all existing attention kernels that the input tokens belong in a consecutive context region in the prefill phase.
Handling non-contiguous KV cache.
@@
Figure 6. Multi-token attention with non-contiguous KV cache. The batch contains two requests, whose tokens (Query representations) are shown on the left. The right side shows the requests’ context (KV-tokens) which reside in non-contiguous GPU memory (block 3, 1 for request 0 and block 0, 4, 5, 2, 7 for request 1).
Existing systems (Yu et al., 2022; Kwon et al., 2023) batch requests separately for the prefill and generation phase so that they can use existing high-performance attention kernels (Dao et al., 2022; Rabe and Staats, 2021) for the prefill phase. Unfortunately, we cannot simply adopt such a design in our setting. This is because existing attention kernels for prefill assume a KV cache with contiguous memory. However, in order to support the swapping of KV-tokens between GPU and CPU, it is more efficient to allow KV-tokens to reside in non-contiguous GPU memory regions. Although vLLM (Kwon et al., 2023) has developed the PagedAttention kernel to handle non-contiguous KV cache, it is designed to be solely used in the generation phase, because it limits each request in the batch to have exactly one input token. As each request has more than one token in the prefill phase, one cannot simply use PagedAttention for prefill. A naive hack is to process the new prompt one token at a time, in the same manner as iterative generation, so that the PagedAttention kernel can be applied. But this method gives up the parallelization opportunity brought by the extra query token dimension in the prefill phase. Thus, to achieve efficient GPU computation, we must address the challenge of supporting non-contiguous KV cache during prefill. Doing so also brings extra benefit: as both the prefill phase and generation phase can compute using the same non-contiguous KV cache, we can handle requests in different phases together in the same batch. Figure 6 illustrates the desired Multi-token attention kernel for computing attention for a batch of two requests, one in its generation phase, the other in its prefill phase, over non-contiguous KV cache.